# Relational Database Distillation: From Structured Tables to Condensed Graph Data

### v1

@@ 1. Introduction
Relational databases (RDBs) are the foundation of global data management systems (Jatana et al., 2012) and estimated to store over 70% of the world’s structured data (DB-Engines, 2023). By organizing information into multiple interlinked tables connected through primary–foreign key relations, RDBs capture both rich structural dependencies and diverse statistical patterns. This generalizable framework serves as the foundational backbone for data storage across critical domains (Ilyas et al., 2008), including finance, education, and e-commerce. For example, in social media platforms (Yin et al., 2015), the relational paradigm models users, posts, comments, and interactions as interdependent entities distributed across multiple tables. This enables a wide range of predictive tasks such as user behavior modeling, recommendation and influence estimation, demonstrating the crucial role of RDBs in understanding large-scale systems.
@@
With the rapid growth of data-driven services, RDBs have expanded to unprecedented scales, often containing billions of rows distributed across interdependent tables. While this growth underscores their central role in data-centric applications, it also introduces significant challenges for predictive modeling (Vogel et al., 2024), particularly in capturing the heterogeneous table attributes and inter-table dependencies induced by foreign-key relations. In an influence prediction task on social media platforms, for example, modeling user influence requires integrating information across multiple tables. The Users table provides profile and demographic attributes, the Posts and Comments tables contain behavioral and content features, and the Interactions table captures the network structure through user identifiers. Effectively combining these heterogeneous attributes and inter-table dependencies is crucial for accurate influence estimation. However, traditional tabular learning pipelines (Fey et al., 2024) rely on extensive feature engineering to join and aggregate related table entities. This process is too labor-intensive to scale and often sacrifices coverage of multi-hop relational information that provides valuable predictive signals.
@@
To alleviate the intricate inter-table dependencies, recent research (Zhang et al., 2023; Robinson et al., 2024; Cvitkovic, 2020; Dwivedi et al., 2025; Chen et al., 2025; Li et al., 2025) has introduced a promising paradigm that reformulates relational databases as heterogeneous graphs (Hu et al., 2020). In this paradigm, table rows/entities are treated as nodes and primary–foreign key relations are modeled as edges. This reformulation enables heterogeneous graph neural networks (GNNs) to capture inter-table dependencies and multi-hop relations through message passing (Gilmer et al., 2017), thereby producing expressive node representations for downstream tasks. Benefiting from graph representation learning, these methods eliminate the need for manual feature engineering and have achieved state-of-the-art performances.
@@
Despite its potential to capture relational information, scaling GNNs to large RDBs remains a formidable challenge. GNNs rely on iterative message passing, which requires repeatedly fetching and aggregating neighboring information across multiple interconnected tables. In large-scale RDBs, this leads to prohibitive training time and high computational cost, compounded by substantial storage overhead. As a result, while GNNs advance predictive performance, their scalability barriers limit practical deployment in resource-constrained or time-sensitive environments.
@@
To address these limitations, we investigate Relational Database Distillation (RDD), a new data-centric paradigm that compresses a large relational database into a compact representation while preserving its essential predictive information. The synthetic data enables much more efficient model training yet achieves performance comparable to training on the original large-scale database. While dataset distillation (Wang et al., 2018) was first explored in computer vision, synthesizing compact data for relational databases introduces several unique challenges due to the nature of tables and relational structures. (i) Multi-modal attributes within tables. Columns in relational tables are often heterogeneous in modality (Cvitkovic, 2020), ranging from numerical and categorical to temporal values, and they are inherently unordered. Consequently, each modality necessitates dedicated encoding and distillation strategies to preserve the representativeness of the compressed data and ensure its utility for model training and downstream tasks. (ii) Heterogeneous tables and complex structural dependencies. RDBs consist of heterogeneous tables linked by foreign-key relations (Fey et al., 2024), where child tables depend on parent tables. This hierarchical dependency compels conventional generation methods (Patki et al., 2016) to construct parent tables before child tables, thereby increasing modeling complexity and incurring substantial generation overhead. Preserving such dependencies in a compressed form is particularly challenging, as it requires capturing both the heterogeneity of table attributes and the inter-table relationships. (iii) Diverse downstream tasks. Due to their extensive industrial adoption, relational databases are expected to support diverse predictive tasks across a broad range of applications (Robinson et al., 2024). While dataset distillation has been explored in modalities such as images (Cazenavette et al., 2022), videos (Wang et al., 2024) and graphs (Gao et al., 2024b), existing methods are predominantly designed for classification and fail to generalize to other tasks (e.g., regression). Therefore, in various downstream tasks, an effective RDD approach is expected to maintain consistent utility of the synthetic data.
@@
  • Through extensive experiments on real-world relational databases, we show that T2G significantly reduces data size while maintaining competitive performances on classification and regression tasks, demonstrating its effectiveness as a scalable paradigm for learning on relational databases.
@@ 2.1. Relational Database
Each entity $v\in T$ is represented as a triplet $v=(p_{v},\mathcal{K}_{v},\textbf{x}_{v})$. Specifically, $p_{v}$ denotes the primary key that uniquely identifies the entity $v$. $\mathcal{K}_{v}\subseteq\{p_{v^{\prime}}:v^{\prime}\in T^{\prime}\wedge(T,T^{\prime})\in\mathcal{L}\}$ represents the set of foreign keys, which define the links from entity $v\in T$ to entities $v^{\prime}\in T^{\prime}$, where $p_{v^{\prime}}$ is the primary key of $v^{\prime}$ in table $T^{\prime}$. An entity can contain one or more foreign keys, depending on the schema design. Finally, $\textbf{x}_{v}$ denotes the multi-modal attributes, which include numerical and categorical features  [^1] We follow previous works (Jurkovic et al., [n. d.]; Ketata et al., 2025) to treat the temporal features as the numerical features.. Formally, $\textbf{x}_{v}$ is a tuple of attribute values: $\textbf{x}_{v}=(\textbf{x}_{v}^{\text{num}},\textbf{x}_{v}^{\text{cat}})$, where $\textbf{x}_{v}^{\text{num}}=(x_{v,1}^{\text{num}},\ldots,x_{v,d_{\text{num}}}^{\text{num}})\in\mathbb{R}^{d_{\text{num}}}$ denotes the numerical attributes. $\textbf{x}_{v}^{\text{cat}}=(x_{v,1}^{\text{cat}},\ldots,x_{v,d_{\text{cat}}}^{\text{cat}})$ denotes the categorical attributes and each $x_{v,i}^{\text{cat}}\in\{1,\ldots,C_{i}\}$ denotes the discrete category index of the $i$-th attribute.
@@ 3.1. Modality-specific Tokenizers
Importantly, each tokenizer can be viewed as a distilled representation of its corresponding column: the original column values serve as keys to retrieve the embeddings from the tokenizer, effectively replacing the raw data with a lightweight yet informative mapping. Thus, storing $\mathcal{M}=\{\mathbf{W}^{\text{num}},\mathbf{E}^{\text{cat}}\}$ captures the essential column-level knowledge of the database, enabling efficient storage, transfer, and inference while preserving task-relevant information. We empirically justify the compact size of the tokenizers through experiments presented in Section 4.6.
@@ 3.5. Further Detailed Analysis
Time complexity. The overall pipeline of T2G comprises three main components: pretraining, graph structure generation, and feature distillation. We analyze the time complexity of each component separately. For the original relational database, let $n_{T}$, $e_{T}$, $d_{{num}}$, and $d_{{cat}}$ denote the average number of entities, links, numerical columns, and categorical columns per table, respectively. The numbers of table and link types are represented by $n$ and $m$. The average number of entities in the synthetic REG is denoted by $n^{\prime}_{T}$, and the token and embedding dimensions are $d_{{token}}$ and $d$.
@@ 4. Experiments
We design comprehensive experiments to validate the efficacy of our proposed methods and explore the following research questions:
Q1: Compared to the other dataset reduction and distillation methods, can T2G achieve better performance?
Q2: Can the synthetic data generated by T2G generalize well to different model architectures?
Q3: How do the different components, i.e., the SBM module, loss terms, and pseudo-labels, influence the performance of T2G?
Q4: How do the hyperparameters (the weight $\beta$ and the sparsity $\rho$) affect the behavior of T2G?
Q5: How effectively does T2G reduce storage size compared to the original database?
@@ 4.1. Experimental Setup
Databases & Baselines. We evaluate our proposed methods on three real-world databases: Rossmann, Walmart, and Airbnb, using the same split method as in the RDBs benchmark (Jurkovic et al., [n. d.]). These databases differ in the types of relationships, as well as in the number of tables and columns. A summary of key characteristics is provided in Table 1.
@@
Table 2. Comparison of downstream model performance trained with different dataset reduction methods. “Whole Database” denotes the model trained on the full original RDB. $r$ denotes the compression ratio. Rossmann and Walmart are regression tasks evaluated by MAE, while Airbnb is a classification task evaluated by AUC (%).
  Database $r$ Baselines Ours
Whole
Database
@@ 4.5. Hyper-parameter Sensitivity Analysis (Q4)
4.6. Storage Efficiency Analysis (Q5)
5. Conclusion
In this paper, we presented Table-to-Graph (T2G), the first framework for relational database distillation, which compresses large-scale relational databases into compact heterogeneous graphs while preserving predictive performance. T2G introduces a clustering-based pretraining objective that enables lightweight, modality-specific tokenization for multi-modal table attributes. Moreover, a SBM mechanism is employed to reconstruct heterogeneous graph structures guided by pseudo-labels. To support both classification and regression tasks, T2G incorporates a KRR-based distillation objective that effectively transfers predictive knowledge into the synthetic features. Extensive experiments on real-world relational databases demonstrate that T2G significantly reduces storage requirements while maintaining competitive performance, establishing a new data-centric paradigm for efficient and scalable learning on relational databases.

### v2

@@ Abstract.
Relational database, Dataset distillation, Multi-tasks, Efficiency.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792734isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Neural networks
@@ 1. Introduction
Relational databases (RDBs) are the foundation of global data management systems (Jatana et al., 2012) and estimated to store over 70% of the world’s structured data (DB-Engines, 2023). By organizing information into multiple interlinked tables connected through primary–foreign key relations, RDBs capture both rich structural dependencies and diverse statistical patterns. This generalizable framework serves as the foundational backbone for data storage across critical domains, including finance, education, and e-commerce (Ilyas et al., 2008; Yin and Cui, 2016; Wang et al., 2025a; Yin et al., 2025; Wang et al., 2024b). For example, in social media platforms (Yin et al., 2015; Hung et al., 2017), the relational paradigm models users, posts, comments, and interactions as interdependent entities distributed across multiple tables. This enables a wide range of predictive tasks such as user behavior modeling, recommendation and influence estimation, demonstrating the crucial role of RDBs in understanding large-scale systems.
@@
With the rapid growth of data-driven services, RDBs have expanded to unprecedented scales, often containing billions of rows distributed across interdependent tables. While this growth underscores their central role in data-centric applications, it also introduces significant challenges for predictive modeling (Vogel et al., 2024; Tran et al., 2025b, a), particularly in capturing the heterogeneous table attributes and inter-table dependencies induced by foreign-key relations. In an influence prediction task on social media platforms, for example, modeling user influence requires integrating information across multiple tables. The Users table provides profile and demographic attributes, the Posts and Comments tables contain behavioral and content features, and the Interactions table captures the network structure through user identifiers. Effectively combining these heterogeneous attributes and inter-table dependencies is crucial for accurate influence estimation. However, traditional tabular learning pipelines (Fey et al., 2024) rely on extensive feature engineering to join and aggregate related table entities. This process is too labor-intensive to scale and often sacrifices coverage of multi-hop relational information that provides valuable predictive signals.
@@
To alleviate the intricate inter-table dependencies, recent research (Zhang et al., 2023; Robinson et al., 2024; Cvitkovic, 2020; Dwivedi et al., 2025; Chen et al., 2025; Li et al., 2025) has introduced a promising paradigm that reformulates relational databases as heterogeneous graphs (Hu et al., 2020). In this paradigm, table rows/entities are treated as nodes and primary–foreign key relations are modeled as edges. This reformulation enables heterogeneous graph neural networks (GNNs) (Liu et al., 2025; Wang et al., 2025b, b; Gao et al., 2023) to capture inter-table dependencies and multi-hop relations through message passing (Gilmer et al., 2017), thereby producing expressive node representations for downstream tasks. Benefiting from graph representation learning, these methods eliminate the need for manual feature engineering and have achieved state-of-the-art performances.
@@
Despite its potential to capture relational information, scaling GNNs to large RDBs remains a formidable challenge. GNNs rely on iterative message passing, which requires repeatedly fetching and aggregating neighboring information across multiple interconnected tables. In large-scale RDBs, this leads to prohibitive training time and high computational cost, compounded by substantial storage overhead. As a result, while GNNs advance predictive performance, their scalability barriers limit practical deployment in resource-constrained or time-sensitive environments (Gao et al., 2024d, c).
@@
To address these limitations, we investigate Relational Database Distillation (RDD), a new data-centric paradigm that compresses a large relational database into a compact representation while preserving its essential predictive information. The synthetic data enables much more efficient model training yet achieves performance comparable to training on the original large-scale database. While dataset distillation (Wang et al., 2018; Gao et al., 2025b, a) was first explored in computer vision, synthesizing compact data for relational databases introduces several unique challenges due to the nature of tables and relational structures (Gao et al., 2025c, 2024b). (i) Multi-modal attributes within tables. Columns in relational tables are often heterogeneous in modality (Cvitkovic, 2020), ranging from numerical and categorical to temporal values, and they are inherently unordered. Consequently, each modality necessitates dedicated encoding and distillation strategies to preserve the representativeness of the compressed data and ensure its utility for model training and downstream tasks. (ii) Heterogeneous tables and complex structural dependencies. RDBs consist of heterogeneous tables linked by foreign-key relations (Fey et al., 2024), where child tables depend on parent tables. This hierarchical dependency compels conventional generation methods (Patki et al., 2016) to construct parent tables before child tables, thereby increasing modeling complexity and incurring substantial generation overhead. Preserving such dependencies in a compressed form is particularly challenging, as it requires capturing both the heterogeneity of table attributes and the inter-table relationships. (iii) Diverse downstream tasks. Due to their extensive industrial adoption, relational databases are expected to support diverse predictive tasks across a broad range of applications (Robinson et al., 2024). While dataset distillation has been explored in modalities such as images (Cazenavette et al., 2022), videos (Wang et al., 2024a) and graphs (Gao et al., 2025d), existing methods are predominantly designed for classification and fail to generalize to other tasks (e.g., regression). Therefore, in various downstream tasks, an effective RDD approach is expected to maintain consistent utility of the synthetic data.
@@
  • Through extensive experiments on real-world relational databases, we show that T2G significantly reduces data size while maintaining competitive performances on classification and regression tasks, demonstrating its effectiveness as a scalable paradigm for learning on relational databases. Our codes are available at: https://github.com/XYGaoG/T2G
@@ 2.1. Relational Database
Each entity $v\in T$ is represented as a triplet $v=(p_{v},\mathcal{K}_{v},\textbf{x}_{v})$. Specifically, $p_{v}$ denotes the primary key that uniquely identifies the entity $v$. $\mathcal{K}_{v}\subseteq\{p_{v^{\prime}}:v^{\prime}\in T^{\prime}\wedge(T,T^{\prime})\in\mathcal{L}\}$ represents the set of foreign keys, which define the links from entity $v\in T$ to entities $v^{\prime}\in T^{\prime}$, where $p_{v^{\prime}}$ is the primary key of $v^{\prime}$ in table $T^{\prime}$. An entity can contain one or more foreign keys, depending on the schema design. Finally, $\textbf{x}_{v}$ denotes the multi-modal attributes, which include numerical and categorical features  [^1] We follow previous works (Jurkovic et al., ; Ketata et al., 2025) to treat the temporal features as the numerical features.. Formally, $\textbf{x}_{v}$ is a tuple of attribute values: $\textbf{x}_{v}=(\textbf{x}_{v}^{\text{num}},\textbf{x}_{v}^{\text{cat}})$, where $\textbf{x}_{v}^{\text{num}}=(x_{v,1}^{\text{num}},\ldots,x_{v,d_{\text{num}}}^{\text{num}})\in\mathbb{R}^{d_{\text{num}}}$ denotes the numerical attributes. $\textbf{x}_{v}^{\text{cat}}=(x_{v,1}^{\text{cat}},\ldots,x_{v,d_{\text{cat}}}^{\text{cat}})$ denotes the categorical attributes and each $x_{v,i}^{\text{cat}}\in\{1,\ldots,C_{i}\}$ denotes the discrete category index of the $i$-th attribute.
@@ 3.1. Modality-specific Tokenizers
Importantly, each tokenizer can be viewed as a distilled representation of its corresponding column: the original column values serve as keys to retrieve the embeddings from the tokenizer, effectively replacing the raw data with a lightweight yet informative mapping. Thus, storing $\mathcal{M}=\{\mathbf{W}^{\text{num}},\mathbf{E}^{\text{cat}}\}$ captures the essential column-level knowledge of the database, enabling efficient storage, transfer, and inference while preserving task-relevant information. We empirically justify the compact size of the tokenizers through experiments presented in Section A.2.
@@ 3.5. Further Detailed Analysis
The detailed time complexity is provided in the Appendix A.1.
@@ 4. Experiments
We design comprehensive experiments to validate the efficacy of our proposed methods and explore the following research questions: Q1: Compared to the other dataset reduction and distillation methods, can T2G achieve better performance? Q2: Can the synthetic data generated by T2G generalize well to different model architectures? Q3: How do the different components, i.e., the SBM module, loss terms, and pseudo-labels, influence the performance of T2G? Q4: How do the hyperparameters (the weight $\beta$ and the sparsity $\rho$) affect the behavior of T2G? Q5: How effectively does T2G reduce storage size compared to the original database?
The results for Q5 are deferred to Appendix A.2.
@@ 4.1. Experimental Setup
Databases & Baselines. We evaluate our proposed methods on 4 real-world databases: Rossmann, Walmart, Airbnb and IMDB, using the same split method as in the RDBs benchmark (Jurkovic et al., ). These databases differ in the types of relationships, as well as in the number of tables and columns. A summary of key characteristics is provided in Table 1.
@@
Table 2. Comparison of model performance trained with different dataset reduction methods. Values are reported as mean ± standard deviation. “Whole Database” denotes the model trained on the full original RDB. $r$ is the compression ratio. Rossmann and Walmart are regression tasks evaluated by MAE, while Airbnb and IMDB are classification tasks evaluated by AUC (%).
  Database $r$ Baselines Ours Whole Database
@@ 4.2. Effectiveness Comparison (Q1)
Our evaluated databases exhibit diverse characteristics. For example, Airbnb contains highly imbalanced categorical columns, with imbalance ratio ranging from 2% to 0.001%. Despite these skewed distributions, our method addresses imbalance through independent tokenizers that assign each category a separate embedding. These embeddings are independently optimized on the original data, preventing majority categories from overwhelming minority representations.
In addition, the relation sparsity of our evaluated databases ranges from 0.1% to 2%. Despite this variation, our SBM module adaptively generates synthetic edges using the cluster-level sparsity defined in Eqs. (7)–(9). By modeling connection probabilities between clusters rather than relying on global sparsity, it captures local interaction patterns and handles diverse sparsity effectively.
@@ 4.5. Hyper-parameter Sensitivity Analysis (Q4)
5. Conclusion
In this paper, we presented Table-to-Graph (T2G), the first framework for relational database distillation, which compresses large-scale relational databases into compact heterogeneous graphs while preserving predictive performance. T2G introduces a clustering-based pretraining objective that enables lightweight, modality-specific tokenization for multi-modal table attributes. Moreover, a SBM mechanism is employed to reconstruct heterogeneous graph structures guided by pseudo-labels. To support both classification and regression tasks, T2G incorporates a KRR-based distillation objective that effectively transfers predictive knowledge into the synthetic features. Extensive experiments on real-world relational databases show that T2G significantly reduces storage requirements while maintaining competitive performance. As an initial attempt at relational database distillation, T2G focuses on a single target table and task, leaving extensions to cross-table and cross-task settings for future work.
Acknowledgements.
The Australian Research Council partially supports this work under the streams of Future Fellowship (Grant No. FT210100624), the Discovery Project (Grant No. DP240101108 and DP260100326), and the Linkage Project (Grant No. LP230200892 and LP240200546).
Appendix A Appendix
A.1. Time Complexity
The overall pipeline of T2G comprises three main components: pretraining, graph structure generation, and feature distillation. We analyze the time complexity of each component separately. For the original relational database, let $n_{T}$, $e_{T}$, $d_{{num}}$, and $d_{{cat}}$ denote the average number of entities, links, numerical columns, and categorical columns per table, respectively. The numbers of table and link types are represented by $n$ and $m$. The average number of entities in the synthetic REG is denoted by $n^{\prime}_{T}$, and the token and embedding dimensions are $d_{{token}}$ and $d$.
A.2. Storage Efficiency Analysis (Q5)